prove that all ellipses with same major axes have the same energy

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Rahul
 
Semi major axis 
a= J²/mc(1 - e²)
1-e² = J²/mca
Now
e² -1 = 2E J²/mc²
Now on solving
a = -c/2E
Hence.E = -c/2a
a constant for the same value of a.
This shows that all ellipses with same major axes have the same energy.

Regards
Arun (askIITians forum expert)